# Scaling MLPs: A Tale of Inductive Bias

### v1

@@ 5 Related Works
There are some prior works that investigate MLPs on vision tasks. Lin et al., (2015) study the performance of MLPs on small scale datasets such as CIFAR10. They observe similar improvements when using inverted bottleneck layers but do not study larger-scale setups, transfer-learning nor do they discuss the implications for theoretical works. Urban et al., (2017) study to what degree convolutions are necessary for good performance and conclude that even with distillation techniques it remains very difficult to train performant MLPs on CIFAR10. Other approaches have focused on sparsifying fully-connected layers through evolutionary training (Mocanu et al.,, 2018; Fernando et al.,, 2016), aiming to learn a good inductive bias from scratch. Similarly, Neyshabur, (2020) study how the inductive bias of MLPs can be improved by systematically sparsifying them with a LASSO-type algorithm, making them more convolution-like. In contrast to these works, our goal is not to enhance the inherent inductive bias of MLPs but study whether it can be overcome with enough scale. The advent of the MLP-Mixer (Tolstikhin et al.,, 2021) has led to a series of follow-up work, similarly using MLPs as a patch processor and token mixer (Touvron et al.,, 2021; Chen et al.,, 2022; Lian et al.,, 2022; Guo et al.,, 2021). Again, we remark that these architectures all possess significantly more inductive bias.
@@ A.5 Computational Efficiency
Appendix B Weight visualizations
@@
Appendix C Inverted Bottleneck MLP Code
We provide PyTorch-style pseudo-code for the inverted bottleneck MLP to highlight its simplicity.
1from torch import nn
2
3class Block(nn.Module):
4    def __init__(self, dim, expansion_factor=4, dropout=0.):
5        super().__init__()
6        self.fn = nn.Sequential(
7            nn.Linear(dim, expansion_factor),
8            nn.GELU(),
9            nn.Dropout(dropout),
10            nn.Linear(expansion_factor, dim),
11            nn.Dropout(dropout)
12        )
13        self.ln = nn.LayerNorm(dim)
14
15    def forward(self, x):
16        return x + self.fn(self.ln(x))
17
18
19def MLP(image_size, channels, dim, depth, num_classes, expansion_factor=4, dropout=0.):
20    return nn.Sequential(
21        nn.Flatten(start_dim=1, end_dim=-1),
22        nn.Linear(image_size * image_size * channels, dim),
23        *[Block(dim, expansion_factor, dropout) for _ in range(depth)],
24        nn.Linear(dim, num_classes)
25    )

### v3

@@ 5 Related Works
There are some prior works that investigate MLPs on vision tasks. Lin et al., (2015) study the performance of MLPs on small scale datasets such as CIFAR10. They observe similar improvements when using inverted bottleneck layers but do not study larger-scale setups, transfer-learning nor do they discuss the implications for theoretical works. The bottleneck structure used in this work has also been investigated theoretically (Parhi and Nowak,, 2021; Shenouda et al.,, 2023; Parkinson et al.,, 2023), further highlighting that such an architecture exhibits desirable properties. Urban et al., (2017) study to what degree convolutions are necessary for good performance and conclude that even with distillation techniques it remains very difficult to train performant MLPs on CIFAR10. Other approaches have focused on sparsifying fully-connected layers through evolutionary training (Mocanu et al.,, 2018; Fernando et al.,, 2016), aiming to learn a good inductive bias from scratch. Similarly, Neyshabur, (2020) study how the inductive bias of MLPs can be improved by systematically sparsifying them with a LASSO-type algorithm, making them more convolution-like. d'Ascoli et al., (2019) on the other hand first train a convolutional network for a certain duration and then subsequently continue training the network as an MLP (by using the correspondence between CNNs and MLPs highlighted in Sec. 2). They show that good performance can be reached if the network was trained long enough as a CNN. In contrast to these works, our goal is not to enhance the inherent inductive bias of MLPs but study whether it can be overcome with enough scale.
The advent of the MLP-Mixer (Tolstikhin et al.,, 2021) has led to a series of follow-up work, similarly using MLPs as a patch processor and token mixer (Touvron et al.,, 2021; Chen et al.,, 2022; Lian et al.,, 2022; Guo et al.,, 2021; Liu et al.,, 2021). Again, we remark that these architectures all possess significantly more inductive bias.
Finally, we would like to remark that MLPs are successfully used in other areas such as novel view synthesis (e.g. NeRF (Mildenhall et al.,, 2021)).
@@ A.2 Additional Ablations
Normalization.
We investigate the importance of the normalization method (LayerNorm vs BatchNorm) in more detail in Table 3. We pre-train two B-MLPs on ImageNet21k with layer normalization and batch normalization and compare the fine-tuning performance on various tasks. We find that the techniques perform similarly, which layer normalisation having a slight edge.
CIFAR-10 CIFAR-100 TinyImageNet ImageNet
LayerNorm $90.0$ $74.6$ $59.6$ $36.2$
BatchNorm $89.4$ $73.8$ $57.7$ $35.9$
Table 3: Pretraining a B-6/Wi-1024 B-MLP with BatchNorm and LayerNorm on ImageNet21k and subsequently fine-tuning.
Label smoothing.
We further ablate the influence of label smoothing on the downstream performance. We pre-train B-MLPs with varying amounts of label smoothing ($\alpha\in\{0.0,0.1,0.3\}$) and evaluate the resulting down-stream fine-tuning performance. We report the results in Table 4. While label smoothing does provide some boost in performance, the gains are very modest. Label smoothing is thus helpful but not essential for training MLPs.
CIFAR10 CIFAR100 TinyImageNet ImageNet
$\alpha=0.3$ $90.0$ $74.6$ $59.6$ $36.2$
$\alpha=0.1$ $89.5$ $73.7$ $58.2$ $36.0$
$\alpha=0.0$ $89.2$ $72.2$ $57.1$ $35.7$
Table 4: Pretraining a B-6/Wi-1024 B-MLP with different amounts of label smoothing on ImageNet21k and subsequently fine-tuning.
@@ A.5 Computational Efficiency
Appendix B Results for Standard MLPs
B.1 Transfer Learning
For completeness we also analyze the transfer performance of standard MLPs when pre-trained on ImageNet21k. We compare a S-MLP of depth $6$ and width $2048$ against a B-MLP of depth $6$ and width $1024$, ensuring that both models roughly have a parameter count of around $\approx 70$ million. We display the results in Table 7. We observe that even the features of a standard MLP (i.e. without residual connections and bottleneck structure) transfer very well on different downstream task. Tthe inverted-bottleneck MLP however still remains superior.
CIFAR10 CIFAR100 TinyImageNet ImageNet
S-MLP $87.1$ $68.3$ $52.1$ $30.2$
B-MLP $90.0$ $74.6$ $59.6$ $36.2$
Table 7: Comparing a S-MLP of width 2048 and depth 6 pre-trained on ImageNet21k, with a B-6/Wi-1024 B-MLP (both models around 70M params) in terms of fine-tuning performance
B.2 Scaling Laws
We also evaluate the scaling law of standard MLPs by training variously sized models on different subsets of ImageNet21k and subsequently linearly probing the features on CIFAR100. The setting is identical to the one described in 4.4. We observe that also standard MLPs exhibit power-law behaviour. The slope ($0.22$ vs $0.25$) and the intercept ($0.18$ vs $0.16$) are however worse when compared against the inverted-bottleneck MLP.
Figure 10: Test error of standard MLPs on CIFAR100 when linearly transferred as a function of PFLOPS, measured according to Eq.(4), on a log-log scale.
Appendix C Weight visualizations
@@
Appendix D Inverted Bottleneck MLP Code
We provide PyTorch-style pseudo-code for the inverted bottleneck MLP to highlight its simplicity. 1from torch import nn 2 3class Block(nn.Module): 4    def __init__(self, dim, expansion_factor=4, dropout=0.): 5        super().__init__() 6        self.fn = nn.Sequential( 7            nn.Linear(dim, int(expansion_factor * dim)), 8            nn.GELU(), 9            nn.Dropout(dropout), 10            nn.Linear(int(expansion_factor * dim), dim), 11            nn.Dropout(dropout) 12        ) 13        self.ln = nn.LayerNorm(dim) 14 15    def forward(self, x): 16        return x + self.fn(self.ln(x)) 17 18 19def MLP(image_size, channels, dim, depth, num_classes, expansion_factor=4, dropout=0.): 20    return nn.Sequential( 21        nn.Flatten(start_dim=1, end_dim=-1), 22        nn.Linear(image_size * image_size * channels, dim), 23        *[Block(dim, expansion_factor, dropout) for _ in range(depth)], 24        nn.Linear(dim, num_classes) 25    )